OpenStudy (anonymous):

sin^-1x+sin^-1y=2pi/3,then cos^-1x+cos^-1y=?

ganeshie8 (ganeshie8):

\[\large \sin^{-1}x + \sin^{-1}y = 2\pi/3\] say, \(\large \sin^{-1}x = A\) \(\large \sin^{-1}y = B\) \[\large A + B = 2\pi/3\]

ganeshie8 (ganeshie8):

\(\large \sin^{-1}x = A \implies \color{red}{x = \sin A} \) \(\large \sin^{-1}y = B \implies \color{red}{y = \sin B} \) so, \[\large \begin{array} \\ \cos^{-1}\color{red}{x} + \cos^{-1}\color{red}{y} &=\cos^{-1}\color{red}{\sin A} + \cos^{-1}\color{red}{\sin B} \\ &=\cos^{-1}\color{red}{\cos(\pi/2-A)} + \cos^{-1}\color{red}{\cos(\pi/2-B)} \\ &=\color{red}{(\pi/2-A)} + \color{red}{(\pi/2-B)} \\ &=\color{red}{\pi-(A+B)} \\ &=\color{red}{\pi-(2\pi/3)} \\ \end{array} \]

ganeshie8 (ganeshie8):

simplify^^

OpenStudy (anonymous):

pi/3

OpenStudy (anonymous):

oh where to study trigonometry

OpenStudy (anonymous):

yes... thanks \(\ddot\smile\)